OpenStudy (anonymous):

Consider the following. y = tan(cos(x)) (a) Write the composite function in the form f(g(x)) by identifying the inner function u = g(x) and the outer function y = f(u). u = g(x) = ? y = f(u) = ? (b) Find the derivative dy/dx. dy/dx = ?

OpenStudy (anonymous):

Well, the inner function is u=cos(x) and the outer function y=tan(u) Because if you plug it in, you get the same thing you started with. The derivative is using the gain rule or: u'f'(u). So u'=-sin(x) and y'=sec^2(u). That gives: -sin(x)sec^2(cos(x))

OpenStudy (anonymous):

u = g(x) = cos x y = f(u) = tanu d/dx cos x = -sinx d/du tan u = sec^2 u then d/du tan(cosx) = sec^2(cosx)(-sin x)

OpenStudy (anonymous):

"...is using the CHAIN rule..."

OpenStudy (anonymous):

g(x)=cosx f(u)=tanu dy/dx=sec2(cosx)*-sinx